iamdead:

A seal dove 95 meters below the surface of the water to reach a fish. It traveled 18% of the total distance every second. What was the depth of the seal after 5 seconds? –17.1 meters –34.2 meters –85.5 meters –68.4 meters

MoonShadow:

What do u think?

MoonShadow:

Ik z answer

Shadow:

They want us to find the seals depth after 5 seconds. First you want to find 18% of 95 (since we knows the seal travels 18% of the total distance of 95, every second). \[95m \times 0.18 = 17.1m/s\] Next, we multiply the rate of 17.1m per second by 5 seconds, in order to get his distance traveled after 5 seconds at that rate. \[\frac{ 17.1m }{ s } \times 5s = 85.5m\] Since they set y = 0 at the surface of the water, traveling downward would mean the values are negative, that means that it would be \[-85.5m\]